The wave function Ψ(x, t) = A exp [i(k1x − ω1t)] + A exp [i(k2x − ω2t)] is a superposition of two free-

particle wave functions. Both k1 and k2 are positive.

(a) Show that this wave function satisfies the Schodinger equation for a free particle of mass m.

(b) Find and plot the probability distribution function for Ψ(x, 0).


1
Expert's answer
2021-05-14T09:42:35-0400

(a)

If particles can exhibit wave nature, they should be described by a function that satisfies a ”wave equation”.


For classical waves


"\dfrac{\partial^2y(x,t)}{\partial x^2}=\dfrac{1}{v^2}\dfrac{\partial ^2y(x,t)}{\partial t^2}"

"f(x,t)=A cos (kx − ωt) + B sin (kx − ωt)\space\space\space\space\space\space\space\space\space\space\space\space ω = vk"


For free-particle waves


"E =\dfrac{p^2}{2m}\space\space \space \space \space \space \space \space \space \space \space \space \space \space \space E = hf = \hbar ω \space \space \space \space \space \space \space \space \space \space \space \space \space p=\dfrac hλ= \hbar k"

"\hbar\omega=\dfrac{\hbar^2k^2}{2m}"


The Schrodinger Equation for a free particle


"-\dfrac{\hbar}{2m}\dfrac{\partial^2\Psi(x,t)}{\partial x^2}=i\hbar\dfrac{\partial\Psi(x,t)}{\partial t}"


The Wave Function


"\u03a8(x, t) = A cos (kx \u2212 \u03c9t) + iA sin (kx \u2212 \u03c9t) = Ae^{i(kx\u2212\u03c9t)}"


(b)

"\Psi(x,t)=A[cos(k_1x-\omega_1t)+sin(k_1x-\omega_1 t)]+A[cos(k_2x-\omega_2t)+sin(k_2x-\omega_2 t)]"

"\Psi(x,t)=2|A|^2(1+cos((k_1-k_2)x-(\omega_1-\omega_2)t))"

"\Psi(x,0)=2|A|^2(1+cos((k_1-k_2)x))"